"Find three consecutive even numbers whose sum is -90"
thanks

Answer by checkley75(3666)

x+(x+2)+(x+4)=-90
3x+6=-90
3x=-90-6
3x=-96
x=-32 answer for the first even number.
-32+2=-30 the second even number
-32+4=-28 the last even number
proof
-32-30-28=-90
-90=-90